OpenStudy (anonymous):

write the equation in slope intercept form of the line that has a slope of (4,-7) and (0,5)

OpenStudy (anonymous):

The answer is y=-3x+5,there are two ways to solve this problem.#1 use the function (y+7)/(5+7)=(x-4)/(0-4) then solve the equation,you'll get the answer.#2 you could calculate the slope of the function which is(5+7)/(0-4)=-3,then we get the function y=-3x+c,plug the (0,5)into the function,we could figure out c equals to 5, also, we could see the value of c from the y-intercept which is 5 as its (0,5) point. Help this answered your question.